# Did we miss P In CAP? Partial Progress Conjecture under Asynchrony

### v1

@@ I Introduction
To illustrate our partial progress conjecture in practice, we deploy our Cassandra protocol in the wild over the Twitter data. We aim to capture the impacts of delays introduced by network partitions on social interactions such as tweeting (with mentions) and re-tweeting along with quoting, replying, and liking a tweet. Our results show that once the network is restored, the system throughput peeks at 6×6$\times$6 × of the steady state due to partial progress made during the partition.
@@ II-B Failure-Free Flow
In Figure 2, we schematically represent our Cassandra protocol. Cassandra requires two phases to achieve consensus among its replicas. The first phase aims to select the strongest proposer (§II-C), while the second phase aims to commit the selected proposal. Cassandra makes use of timers in the first phase for rapid convergence (§II-D); Cassandra is safe and live without timers. Each replica maintains an ordered log of transactions (denoted as $hist$). Next, we describe Cassandra in detail and present its pseudocode in Figure 3.
Proposal. The first phase of Cassandra starts when the $i$-th replica $R_{i}$ has a client transaction ${T}$ that it wants to propose (Line 1). Next, $R_{i}$ creates a proposal $\phi_{i}$, which it broadcasts to all the replicas (Lines 2-5). This proposal includes (1) hash of the ordered log $hist$, (2) client transaction and its hash, and (3) a function $\mathop{\textsf{DRF}}()$ that assigns a rank to $R_{i}$’s proposal (§II-C). Post this, $R_{i}$ sets its proposal as the strongest proposal ($Max$) and waits on a timer ($\tau_{p}$) to receive all the other proposals.
Proposal Selection. While waiting on the timer $\tau_{p}$, replica $R_{i}$ may receive proposals (Line 8) from several replicas. It uses these proposals to determine the strongest proposer. It compares the ordered log ($hist$) of each incoming proposal against $Max$ and sets $Max$ to the stronger proposal if any. As described above, if two proposals have same ordered logs, then $R_{i}$ updates $Max$ to the proposal with higher rank (Line 12-14). Instead, if a received proposal’s log has larger suffix, we update $Max$ to that proposal.
@@
Commit. When a replica $R_{i}$ receives a Prepare message from the $j$-th replica $R_{j}$, it acknowledges this message by sending VotePrep message. Once the replica $R_{j}$ receives identical VotePrep message from $\mathbf{f}{}+1$ replicas, it adds ${T}$ to its log $hist$ and broadcasts Commit message. This Commit message allows all the replicas to eventually commit the transaction.
@@
// $\mathop{\textsf{DRF}}(i)$ determines strength of $i$-th replica’s proposal.// $hist$ is the ordered log of transactions.// $Max$ is the strongest proposal for a round.
@@
1:event Received a client transaction ${T}$ to propose do
@@
16:    Send $\textsc{Vote}(i,\Delta_{Max})$ to replica with $Max$ proposal.
@@ III-A Merge Procedure
    If suffix of $hist_{i}$ is $\{{T}_{x},....,Prep({T}_{y}),...\}$ and suffix of $hist_{j}$ is $\{{T}_{x},...,Prep({T}_{z},...\})$, such that $x<y\wedge x<z$, then $Max$ is set to the replica with the strongest DRF.
@@ III-A1 Protocol Steps
Vote Timeout. If a replica $R_{i}$’s vote timer timeouts while waiting for the Prepare message, it assumes that the replica it voted as the strongest proposer has failed. As a result, it abandons this round (say $r$) of consensus and ignores any proposals from other replicas for round $r$. As $R_{i}$ was never selected as the strongest proposer in round $r$, it adds its own proposal for round $r$ to its log $hist$ and starts a new round of consensus ($r+1$). Notice that $R_{i}$ marks its proposal as unprepared in the log. In round $r+1$, once $R_{i}$ has a new transaction to propose, it constructs a new proposal that extends its log $hist$ and follows the remaining steps in Figure 3.
Prepare Timeout. If a replica $R_{i}$’s prepare timer timeouts while waiting for the Commit message, it assumes that the strongest proposer has failed after sending a Prepare message. As a result, $R_{i}$ adds this strongest proposal to its log and marks its as prepared in its log. Next, $R_{i}$ terminates this round (say $r$) of consensus. When $R_{i}$ has a new transaction to propose, it constructs a new proposal that extends its log $hist$ and follows the remaining steps in Figure 3 to start a new round of consensus.
Conflict Resolution. During the proposal selection phase (§II-B) a replica $R_{i}$ may receive a conflicting proposal $\phi_{j}$ from a replica $R_{j}$. This conflict is the result of $R_{i}$/$R_{j}$ preparing distinct transactions. Fortunately, our rules for determining the strongest proposal dictate which replica’s proposal should be considered strongest. If $R_{i}$’s proposal is weaker, it rollbacks its state.
@@ IV Evaluation
[…]
Promise of Partial Progress. We evaluate the throughput and latency of the Cassandra in partitioned setting where no partition has the majority of replicas. We observe during the partition, Cassandra enters a speculative ordering and execution phase, where transactions are only softly committed and finality is only established once the partitions are recovered. Therefore, once the network is restored, we observe throughout peeks at 6×6$\times$6 × of the steady state made possible by enabling partial progress. As expected, the finality latency is delayed proportionally to the length of the network outage. The results are shown in Figure 4.

### v2

@@ I Introduction
To illustrate our partial progress conjecture in practice, we deploy our Cassandra protocol in the wild over the Twitter data. We aim to capture the impacts of delays introduced by network partitions on social interactions such as tweeting (with mentions) and re-tweeting along with quoting, replying, and liking a tweet. Our results show that once the network is restored, the system throughput peeks at $6\times$ of the steady state due to partial progress made during the partition.
@@ II-B Failure-Free Flow
In Figure 2, we schematically represent our Cassandra protocol. Cassandra requires two phases to achieve consensus among its replicas. The first phase aims to select the strongest proposer (§II-C), while the second phase aims to commit the selected proposal. Cassandra makes use of timers in the first phase for rapid convergence (§II-D); Cassandra is safe and live without timers. Each replica maintains an ordered log of transactions (denoted as histhist). Next, we describe Cassandra in detail and present its pseudocode in Figure 3.
Proposal. The first phase of Cassandra starts when the $i$-th replica $R_{i}$ has a client transaction T{T} that it wants to propose (Line 1). Next, $R_{i}$ creates a proposal $\phi_{i}$, which it broadcasts to all the replicas (Lines 2-5). This proposal includes (1) hash of the ordered log histhist, (2) client transaction and its hash, and (3) a function $\mathop{\textsf{DRF}}()$ that assigns a rank to $R_{i}$’s proposal (§II-C). Post this, $R_{i}$ sets its proposal as the strongest proposal ($Max$) and waits on a timer ($\tau_{p}$) to receive all the other proposals.
Proposal Selection. While waiting on the timer $\tau_{p}$, replica $R_{i}$ may receive proposals (Line 8) from several replicas. It uses these proposals to determine the strongest proposer. It compares the ordered log (histhist) of each incoming proposal against $Max$ and sets $Max$ to the stronger proposal if any. As described above, if two proposals have same ordered logs, then $R_{i}$ updates $Max$ to the proposal with higher rank (Line 12-14). Instead, if a received proposal’s log has larger suffix, we update $Max$ to that proposal.
@@
Commit. When a replica $R_{i}$ receives a Prepare message from the $j$-th replica $R_{j}$, it acknowledges this message by sending VotePrep message. Once the replica $R_{j}$ receives identical VotePrep message from $\mathbf{f}{}+1$ replicas, it adds T{T} to its log histhist and broadcasts Commit message. This Commit message allows all the replicas to eventually commit the transaction.
@@
// $\mathop{\textsf{DRF}}(i)$ determines strength of $i$-th replica’s proposal.// histhist is the ordered log of transactions.// $Max$ is the strongest proposal for a round.
@@
1:event Received a client transaction T{T} to propose do
@@
16:  Send $\textsc{Vote}(i,\Delta_{Max})$ to replica with MaxMax proposal.
@@ III-A Merge Procedure
[…]
@@ III-A1 Protocol Steps
Vote Timeout. If a replica $R_{i}$’s vote timer timeouts while waiting for the Prepare message, it assumes that the replica it voted as the strongest proposer has failed. As a result, it abandons this round (say $r$) of consensus and ignores any proposals from other replicas for round $r$. As $R_{i}$ was never selected as the strongest proposer in round $r$, it adds its own proposal for round $r$ to its log histhist and starts a new round of consensus ($r+1$). Notice that $R_{i}$ marks its proposal as unprepared in the log. In round $r+1$, once $R_{i}$ has a new transaction to propose, it constructs a new proposal that extends its log histhist and follows the remaining steps in Figure 3.
Prepare Timeout. If a replica $R_{i}$’s prepare timer timeouts while waiting for the Commit message, it assumes that the strongest proposer has failed after sending a Prepare message. As a result, $R_{i}$ adds this strongest proposal to its log and marks its as prepared in its log. Next, $R_{i}$ terminates this round (say $r$) of consensus. When $R_{i}$ has a new transaction to propose, it constructs a new proposal that extends its log histhist and follows the remaining steps in Figure 3 to start a new round of consensus.
Conflict Resolution. During the proposal selection phase (§II-B) a replica $R_{i}$ may receive a conflicting proposal $\phi_{j}$ from a replica $R_{j}$. This conflict is the result of $R_{i}$$R_{j}$ preparing distinct transactions. Fortunately, our rules for determining the strongest proposal dictate which replica’s proposal should be considered strongest. If $R_{i}$’s proposal is weaker, it rollbacks its state.
@@ IV Evaluation
Setup and Benchmark. We run our experiments on a local cluster and deploy each replica on a virtual machine having a $4$-core CPU and $8\text{\,}\mathrm{GB}\text{/}$ memory. We run Cassandra on $100$ replicated machines, where the primary mode of communication is gossip to exhibit decentralized deployment. We inject 10,000 transactional tweets per second, where each transaction has a payload of under 1KB. The experiments are run for 260 seconds, where at the 60 seconds mark, the network undergoes partition, the partition lasts for 60 seconds, and the network is restored at 120 seconds. To collect results after reaching a steady state, we discard the measurement during the warmup period, and measurement results are collected over three runs.
Promise of Partial Progress. We evaluate the throughput and latency of the Cassandra in partitioned setting where no partition has the majority of replicas. We observe during the partition, Cassandra enters a speculative ordering and execution phase, where transactions are only softly committed and finality is only established once the partitions are recovered. Therefore, once the network is restored, we observe throughout peeks at $6\times$ of the steady state made possible by enabling partial progress. As expected, the finality latency is delayed proportionally to the length of the network outage. The results are shown in Figure 4.